Determine the market price of a 12-year, 5.15%, $150,000 coupon bond when the market interest rate is 5.75%.

    Note: you must use the pricing formula from class. Do not use 12 different terms.
Answer #1

Coupon payment = 0.0515 * 150000 = 7725

Price of bond = 7725 * (P/A, 5.75%,12) + 150000 * (P/F, 5.75%,12)

= 7725 * 8.499956 + 150000 * 0.511253

= 142350.04